# Metric Rarity and the Rise of Symmetry in Invariant Optimization

### v1

@@ 4.2 Generalization to Finite Group Actions
Theorem 4.5 (Real Image Volume Ratio (W. Sawin)).
@@
A point $y\in Y(\mathbb{R})$ is real if and only if $y=\overline{y}$. Lifting this to $X$, if $\pi(x)=y$ then $\overline{x}=\sigma x$ for some $\sigma\in G$ satisfying $\sigma^{2}=\mathrm{id}$. Define the twisted real subspaces:
@@
Comparison with the Gaussian case validates the use of our induced geometric measure. The Gaussian measure, which represents a standard ”unbiased” distribution on the coefficient space, under Assumption 4.1 dictates an extremely rapid super-exponential decay for the volume of the real image ($\sim e^{-Cn^{2}}$). In contrast, our geometric measure yields a significantly slower decay ($\sim e^{-Cn\log n}$). This suggests that the geometric measure does not artificially suppress the size of the real image; in fact, it attributes to it a relative volume significantly larger than the standard probabilistic baseline.
@@ 4.3 Iteration Along Fixed-Point Strata
This result establishes a monotonicity of ”probability” across strata. Typically, as the stabilizer $H$ increases in size, the effective group $\overline{G}=N_{G}(H)/H$ becomes smaller. Consequently, the number of involutions $\mathrm{Inv}(\overline{G})$ decreases. Since the relative volume of the real image is proportional to $1/\#\mathrm{Inv}(\overline{G})$, the real image becomes less rare in strata associated with higher symmetries. This provides a geometric justification for the recursive emergence of high symmetry: a generic critical point is statistically unlikely to reside in strata where $\overline{G}$ is large (due to the rarity of the real image), and is thus filtered toward strata where $\overline{G}$ is small.
@@ 5 The Physical Setup: Particle Systems and Shape Spaces
Let $q:\mathcal{V}\to X_{\operatorname{shape}}$ denote the quotient map. Crucially, strictly real configurations map to a subset of the real locus of the quotient. We define the Physical Shape Space as this image:
@@ 5.1 Metric Rarity in the Shape Space Quotient
There exists a unique smooth Riemannian metric $g_{X}$ on $X_{\operatorname{shape}}^{\mathrm{\operatorname{pr}}}$ such that the quotient map $q:\mathcal{V}^{{\mathrm{\operatorname{pr}}}}\to X_{\operatorname{shape}}^{{\mathrm{\operatorname{pr}}}}$ restricts to a Riemannian submersion. Crucially, the intrinsic geodesic distance induced by $g_{X}$ coincides exactly with the optimal alignment distance $d_{X}$ defined above.
@@
Crucially, the metric weight of these sectors depends on the geometry of the defining cocycles. With respect to a natural invariant measure, the volume of a sector $Y_{\alpha}$ is proportional to the volume of the conjugacy class associated with $\alpha$ within the maximal compact subgroup $K\subset G$. Our result illustrates two distinct regimes of this principle:
@@ 6 Unified Algebraic Interpretation of Empirical Symmetry
The ubiquitous symmetry observed in [3] and [1, 2] provides a comprehensive empirical validation of ”Regime I”. In these experimental settings, the optimization landscape is governed by the metric rarity of the real image $L\subset Y(\mathbb{R})$. As established in Section 1.2.1, the ”interior” of the quotient space (corresponding to asymmetric configurations) occupies negligible volume. Consequently, generic critical points are statistically forced onto the boundary strata $\partial L$, necessitating non-trivial stabilizers.
@@
The empirical observation that $I_{E}$ remains non-trivial even when $I_{V}$ vanishes confirms our geometric rarity hypothesis. Even if a potential (e.g., $1/r^{2}$) is singular at the collision locus—preventing critical points on the boundary of $L\subset(\mathcal{V}/S_{n})(\mathbb{R})$ (where $I_{V}\neq\{e\}$)—the system settles on the boundary of the shape space $L\subset Y_{\operatorname{shape}}(\mathbb{R})$. The metric rarity of the real image in the final quotient $Y_{\operatorname{shape}}=X_{\operatorname{shape}}\mathbin{/\mkern-6.0mu/}S_{n}$ (Theorem 5.13) is sufficient to force the shape itself to be symmetric ($I_{E}\neq\{e\}$). Moreover, the additional volume factor of $1/2^{\min(d,n)}$ in Theorem 5.13 explains the consistent empirical strict inclusion $I_{V}(x)<I_{E}(x)$ reported in [3].
@@ 7 Conclusion and Future Work
Throughout this work, we have adopted the following general setting: an irreducible complex affine algebraic variety $X$ defined over $\mathbb{R}$, equipped with a faithful action of a finite group $G$ by real automorphisms, and a $G$-invariant regular function $f\in\mathbb{R}[X]^{G}$ viewed as an energy landscape on the real locus $X(\mathbb{R})$. The guiding question was when critical points—and in particular global minimizers—exhibit a non-trivial stabilizer. Across the families of models considered here, we observed two recurring empirical regimes: in Regime I almost all critical points exhibit a non-trivial stabilizer, while in Regime II, a distinct energetic ordering emerges: high-symmetry configurations are systematically concentrated at the bottom of the landscape, with the deepest minima exhibiting significantly larger stabilizer subgroups than the bulk.
@@
For Regime II, we proposed the “Active Constraint” hypothesis. We posit that even when the descended potential $\tilde{f}$ is sufficiently complex to exhibit a rugged landscape within the interior $L^{\circ}$, the global minimum is driven to the boundary $\partial L$ by a coherent global gradient of $\tilde{f}$ that does not vanish in the interior. We found strong empirical support for this mechanism in Lennard-Jones clusters, where the well-documented “funnel” topology effectively directs the system toward a highly symmetric global minimum. In our geometric interpretation, this implies that the global gradient on the quotient space channels the trajectory specifically into a sharp corners of the boundary $\partial L$.
@@
  • The Ambient Landscape Topology. A compelling direction for future inquiry is to understand to what extent the “funnel” topology characteristic of Lennard-Jones clusters is an intrinsic feature of the algebraic potential $\tilde{f}$ on the entire ambient space $Y(\mathbb{R})$. We acknowledge that since $\tilde{f}$ is unbounded below on the ambient quotient (as shown in Remark 5.15), formulating this problem rigorously requires imposing appropriate bounds or regularization. Nevertheless, the fundamental question remains: if one were to analyze the landscape on this extended domain, ignoring the specific geometric constraints of the real image, would the funnel structure persist? Or would the landscape degenerate into a “glassy” structure characterized by frustration and a proliferation of competing minima?
@@
  • Beyond Group Actions: General Morphisms and Deep Learning. The prevalence of symmetry in neural networks has been rigorously documented in what we term Regime I. Arjevani et al. [1] showed that loss functions of shallow ReLU networks systematically preserve the permutation symmetries of hidden neurons, rendering symmetry ubiquitous rather than rare. Furthermore, empirical analyses of machine learning energy landscapes [27] reveal a “funnel-like” topography strikingly similar to the single-funnel landscapes of physical systems like Lennard-Jones clusters, which facilitates efficient relaxation to low-energy states. To provide a unified geometric explanation for these phenomena in Deep Learning, we propose looking beyond group quotients. The phenomenon of real-image compression is not specific to quotients; for a general subjective finite morphism $\phi:X\to Y$ defined over $\mathbb{R}$, we expect the 2-primary part of the degree $\deg(\phi)$ to control the extent to which the real image $\phi(X(\mathbb{R}))$ is “squeezed” inside $Y(\mathbb{R})$. This suggests a speculative geometric lens for analyzing neural network loss landscapes: if a loss function on weight space factors (even approximately) as $f=\tilde{f}\circ\phi$, where $\phi$ quotients out built-in symmetries and structural dependencies, then optimization is effectively constrained to the real image $L_{\phi}:=\phi(X(\mathbb{R}))$. Successive “folds” of $L_{\phi}$ could then offer a geometric route to “deep singularities,” constraining the feasible real locus and pushing global minima toward increasingly singular corners, in direct analogy with the Lennard-Jones phenomenon studied here.
@@ Appendix C Disentangling the Driver of the Active Constraint
One hypothesis, translating the statistical arguments of Wales [4] into our geometric framework, is that this gradient is entropic in nature, driven by the variance profile of the ensemble. Alternatively, the gradient may be a generic topological consequence of optimizing a rugged function over a restricted domain—much like the ”Carpet on the Himalayas” analogy—persisting even in the absence of variance fluctuations.
@@ 2. The Rigorous Precedent: The Bos-Ware Limit
This hypothesis is strongly supported in the one-dimensional case by Bos and Ware [28]. They analyzed the global minimum of random Kac polynomials $p(t)$ on a bounded interval [a,a][-a,a]. They proved that despite the landscape becoming increasingly rugged as the degree $n$ rises, the probability that the global minimum occurs strictly on the boundary (the endpoints $\{-a,a\}$) remains uniformly bounded away from zero:
@@
Crucially, for the ensemble they analyzed, the variance $\sigma^{2}(x)$ is strictly convex and maximized at these endpoints. Thus, in the 1D limit, the boundary preference aligns perfectly with the direction of maximal variance growth.
@@ 3. The Sphere Control Experiment: Testing the Hypothesis
    Crucially, moving along this standard circle in configuration space induces radial movement in the quotient space. Specifically, the quotient norm $\|\pi(x)\|$ varies along $\mathcal{M}_{X\text{-sphere}}$ and attains its maximum exactly at the symmetric configurations (where $x_{1}=x_{2}$). Thus, restricting to the standard sphere $\mathcal{M}_{X\text{-sphere}}$ reintroduces a variance gradient that peaks at the boundary.
@@
    Intrinsic Driver: For complex domains (as in $n=3$), the boundedness and shape of the real image alone are sufficient to push the global minimum to the boundary. The rugged landscape naturally ”slides” into the intricate corners of the domain.

### v2

@@ 4.2 Generalization to Finite Group Actions
Theorem 4.5 (Real Image Volume Ratio).
@@
The proof follows a geometric argument communicated by W. Sawin [26]. A point $y\in Y(\mathbb{R})$ is real if and only if $y=\overline{y}$. Lifting this to $X$, if $\pi(x)=y$ then $\overline{x}=\sigma x$ for some $\sigma\in G$ satisfying $\sigma^{2}=\mathrm{id}$. Define the twisted real subspaces:
@@
Comparison with the Gaussian case validates the use of our induced geometric measure. The Gaussian measure, which represents a standard ”unbiased” distribution on the coefficient space, under Assumption 4.1 dictates an extremely rapid super-exponential decay for the volume of the real image ($\sim e^{-Cn^{2}}$). In contrast, our geometric measure yields a slower decay ($\sim e^{-Cn\log n}$). This suggests that the geometric measure does not artificially suppress the size of the real image; in fact, it attributes to it a relative volume strictly larger than the standard probabilistic baseline.
@@ 4.3 Iteration Along Fixed-Point Strata
This result establishes a monotonicity of ”probability” across strata. Typically, as the stabilizer $H$ increases in size, the effective group $\overline{G}=N_{G}(H)/H$ becomes smaller. Consequently, the number of involutions $\mathrm{Inv}(\overline{G})$ decreases. Since the relative volume of the real image is proportional to $1/\#\mathrm{Inv}(\overline{G})$, the real image becomes less rare in strata associated with higher symmetries. This provides a geometric justification for the recursive emergence of high symmetry: a generic critical point is statistically unlikely to reside in strata where $\overline{G}$ is large (due to the rarity of the real image), and is thus constrained to strata where $\overline{G}$ is small.
@@ 5 The Physical Setup: Particle Systems and Shape Spaces
Let $q:\mathcal{V}\to X_{\operatorname{shape}}$ denote the quotient map. Strictly real configurations map to a subset of the real locus of the quotient. We define the Physical Shape Space as this image:
@@ 5.1 Metric Rarity in the Shape Space Quotient
There exists a unique smooth Riemannian metric $g_{X}$ on $X_{\operatorname{shape}}^{\mathrm{\operatorname{pr}}}$ such that the quotient map $q:\mathcal{V}^{{\mathrm{\operatorname{pr}}}}\to X_{\operatorname{shape}}^{{\mathrm{\operatorname{pr}}}}$ restricts to a Riemannian submersion. The intrinsic geodesic distance induced by $g_{X}$ coincides exactly with the optimal alignment distance $d_{X}$ defined above.
@@
The metric weight of these sectors depends on the geometry of the defining cocycles. With respect to a natural invariant measure, the volume of a sector $Y_{\alpha}$ is proportional to the volume of the conjugacy class associated with $\alpha$ within the maximal compact subgroup $K\subset G$. Our result illustrates two distinct regimes of this principle:
@@ 6 Unified Algebraic Interpretation of Empirical Symmetry
The prevalent symmetry observed in [3] and [1, 2] provides a comprehensive empirical validation of “Regime I”. In these experimental settings, the optimization landscape is governed by the metric rarity of the real image $L\subset Y(\mathbb{R})$. As established in Section 1.2.1, the “interior” of the quotient space (corresponding to asymmetric configurations) occupies negligible volume. Consequently, generic critical points are statistically constrained to the boundary strata $\partial L$, necessitating non-trivial stabilizers.
@@
The empirical observation that $I_{E}$ remains non-trivial even when $I_{V}$ vanishes confirms our geometric rarity hypothesis. Even if a potential (e.g., $1/r^{2}$) is singular at the collision locus—preventing critical points on the boundary of $L\subset(\mathcal{V}/S_{n})(\mathbb{R})$ (where $I_{V}\neq\{e\}$)—the system settles on the boundary of the shape space $L\subset Y_{\operatorname{shape}}(\mathbb{R})$. The metric rarity of the real image in the final quotient $Y_{\operatorname{shape}}=X_{\operatorname{shape}}\mathbin{/\mkern-6.0mu/}S_{n}$ (Theorem 5.13) is sufficient to statistically favor symmetric shapes with non trivial symmetry ($I_{E}\neq\{e\}$). Moreover, the additional volume factor of $1/2^{\min(d,n)}$ in Theorem 5.13 explains the consistent empirical strict inclusion $I_{V}(x)<I_{E}(x)$ reported in [3].
@@ 7 Conclusion and Future Work
Throughout this work, we have adopted the following general setting: an irreducible complex affine algebraic variety $X$ defined over $\mathbb{R}$, equipped with a faithful action of a finite group $G$ by real automorphisms, and a $G$-invariant regular function $f\in\mathbb{R}[X]^{G}$ viewed as an energy landscape on the real locus $X(\mathbb{R})$. The central problem addressed herein is to determine conditions under which critical points—and in particular global minimizers—exhibit a non-trivial stabilizer. Across the families of models considered here, we observed two recurring empirical regimes: in Regime I almost all critical points exhibit a non-trivial stabilizer, while in Regime II, a distinct energetic ordering emerges: high-symmetry configurations are systematically concentrated at the bottom of the landscape, with the deepest minima exhibiting larger stabilizer subgroups than the bulk.
@@
For Regime II, we proposed the “Active Constraint” hypothesis. We posit that even when the descended potential $\tilde{f}$ is sufficiently complex to exhibit a multimodal landscape within the interior $L^{\circ}$, the global minimum appears to be driven to the boundary $\partial L$ by a coherent global gradient of $\tilde{f}$ that does not vanish in the interior. We found strong empirical support for this mechanism in Lennard-Jones clusters, where the well-documented “funnel” topology effectively directs the system toward a highly symmetric global minimum. In our geometric interpretation, this implies that the global gradient on the quotient space channels the trajectory specifically into singular strata of high codimension of the boundary $\partial L$.
@@
  • The Ambient Landscape Topology. A natural question is to what extent the “funnel” topology characteristic of Lennard-Jones clusters is an intrinsic feature of the algebraic potential $\tilde{f}$ on the entire ambient space $Y(\mathbb{R})$. We acknowledge that since $\tilde{f}$ is unbounded below on the ambient quotient (as shown in Remark 5.15), formulating this problem rigorously requires imposing appropriate bounds or regularization. Nevertheless, the fundamental question remains: if one were to analyze the landscape on this extended domain, ignoring the specific geometric constraints of the real image, would the funnel structure persist? Or would the landscape degenerate into a “glassy” structure characterized by frustration and a large number of competing minima?
@@
  • Beyond Group Actions: General Morphisms and Deep Learning. The prevalence of symmetry in neural networks has been rigorously documented in what we term Regime I. Arjevani et al. [1] showed that loss functions of shallow ReLU networks systematically preserve the permutation symmetries of hidden neurons, rendering symmetry prevalent. Empirical analyses of machine learning energy landscapes [31] reveal a “funnel-like” topography analogous to the single-funnel landscapes of physical systems like Lennard-Jones clusters, which facilitates efficient relaxation to low-energy states. To provide a unified geometric explanation for these phenomena in Deep Learning, we propose looking beyond group quotients. The phenomenon of real-image compression is not specific to quotients; for a general subjective finite morphism $\phi:X\to Y$ defined over $\mathbb{R}$, we expect the 2-primary part of the degree $\deg(\phi)$ to control the extent to which the real image $\phi(X(\mathbb{R}))$ is “squeezed” inside $Y(\mathbb{R})$. This suggests a speculative geometric lens for analyzing neural network loss landscapes: if a loss function on weight space factors (even approximately) as $f=\tilde{f}\circ\phi$, where $\phi$ quotients out built-in symmetries and structural dependencies, then optimization is effectively constrained to the real image $L_{\phi}:=\phi(X(\mathbb{R}))$. Successive “folds” of $L_{\phi}$ could then offer a geometric route to “deep singularities,” constraining the feasible real locus and pushing global minima toward increasingly singular corners, in direct analogy with the Lennard-Jones phenomenon studied here.
@@ Appendix B Additional Experiments: Robustness of the Regime II Phenomenon
Numerical tolerances (convergence and symmetry).
Unless stated otherwise, throughout Appendix B (and in the main Figure 1(a)) we use the following fixed tolerances. First, we declare that a Newton trajectory has converged to a (numerical) critical point when the gradient norm falls below
$\|\nabla f(x)\|<\varepsilon,\qquad\varepsilon=10^{-1}.$
For constrained runs on a manifold (e.g. $\|x\|=1$ or $\|e(x)\|=1$), we apply the same rule to the tangent-projected gradient, i.e. $\|\nabla_{T}f(x)\|<\varepsilon$. Second, to compute the symmetry score $D(x)$, we treat two coordinates as equal when $|x_{i}-x_{j}|\leq\delta$ with
$\delta=10^{-2}.$
Operationally, we sort the coordinates of $x$ and set $D(x)=1+\#\{i:\,x_{(i+1)}-x_{(i)}>\delta\}$. The same tolerance $\delta$ is also used when deduplicating numerically detected critical points up to permutation symmetry.
@@ Appendix C Disentangling the Driver of the Active Constraint
One hypothesis, translating the statistical arguments of Wales [4] into our geometric framework, is that this gradient is entropic in nature, driven by the variance profile of the ensemble. Alternatively, the gradient may be a generic topological consequence of optimizing a multimodal function over a restricted domain—much like the “Carpet on the Himalayas” analogy—persisting even in the absence of variance fluctuations.
@@ 2. The Rigorous Precedent: The Bos-Ware Limit
This hypothesis is strongly supported in the one-dimensional case by Bos and Ware [32]. They analyzed the global minimum of random Kac polynomials $p(t)$ on a bounded interval [a,a][-a,a]. They proved that despite the landscape becoming increasingly multimodal as the degree $n$ rises, the probability that the global minimum occurs strictly on the boundary (the endpoints $\{-a,a\}$) remains uniformly bounded away from zero:
@@
For the ensemble they analyzed, the variance $\sigma^{2}(x)$ is strictly convex and maximized at these endpoints. Thus, in the 1D limit, the boundary preference aligns perfectly with the direction of maximal variance growth.
@@ 3. The Sphere Control Experiment: Testing the Hypothesis
    Moving along this standard circle in configuration space induces radial movement in the quotient space. Specifically, the quotient norm $\|\pi(x)\|$ varies along $\mathcal{M}_{X\text{-sphere}}$ and attains its maximum exactly at the symmetric configurations (where $x_{1}=x_{2}$). Thus, restricting to the standard sphere $\mathcal{M}_{X\text{-sphere}}$ reintroduces a variance gradient that peaks at the boundary.
@@
    Intrinsic Driver: For complex domains (as in $n=3$), the boundedness and shape of the real image alone are sufficient to push the global minimum to the boundary. The multimodal landscape naturally attains its minimum at the boundary of the domain.